OpenStudy (thatcreepyboy):

Titans crime is at foot! Find the diameter of a cone that has a volume of 83.74 cubic inches and a height of 5 inches. Use 3.14 for pi. 3 inches 4 inches 8 inches 16 inches

OpenStudy (thatcreepyboy):

lets fight this question and get some pizza!

OpenStudy (jango_in_dtown):

Let the radius be r... Then V=(1/3) *3.14*r^2*h now put the values of V and h

OpenStudy (thatcreepyboy):

r means ratio?

OpenStudy (jango_in_dtown):

so r^2=3V/(3.14*h)=3*83.74/(3.14*5)=16.001 so r =4 i.e. d=2r=8

OpenStudy (jango_in_dtown):

r is radius and d is diameter

OpenStudy (jango_in_dtown):

8inches is the answer

OpenStudy (thatcreepyboy):

Thank you, we shall celebrate with pizza! At the tower!

OpenStudy (thatcreepyboy):

@jango_IN_DTOWN Can you show me how to do this? Find the area of a circle with a diameter of 12 inches. Use 3.14 for pi. 37.68 in2 113.04 in2 339.12 in2 452.16 in2

OpenStudy (jango_in_dtown):

First tell the the formula of the area of a circle

OpenStudy (thatcreepyboy):

uh

OpenStudy (thatcreepyboy):

A = π × r2?

OpenStudy (thatcreepyboy):

NOWAIT tt

OpenStudy (jango_in_dtown):

correct

OpenStudy (jango_in_dtown):

if r is the radius then area= pi * r^2 for a circle

OpenStudy (jango_in_dtown):

now diameter=2*radius

OpenStudy (jango_in_dtown):

so r=d/2 here d=12, so r=6

OpenStudy (jango_in_dtown):

area= pi * 6^2=3.14*36=113.04

OpenStudy (thatcreepyboy):

So you're suppose to divide the numbers by pie?

OpenStudy (jango_in_dtown):

diameter is given.. to get the radius, divide the diameter by 2

OpenStudy (jango_in_dtown):

then area =pi*r^2 here pi=3.14 and r=6

OpenStudy (thatcreepyboy):

Ah okay thank you